A closely wound 0.20m × .20 m square coil of 100 turns is exposed to an external magnetic field of 0.74 T that is perpendicular to the plane of the coil. The coil is then rotated for 10 s until it makes 60 degrees with the magnetic field. Find the average induced electromotive force during the 10 s interval.


1
Expert's answer
2022-04-18T17:31:21-0400

The emf induced in the coil is given by the Faraday's law:


"\mathcal{E} = -\dfrac{\Delta\Phi}{\Delta t}"

where "\Delta\Phi" is the change in magnetic flux, and "\Delta t = 10s" is the time elapsed. The initial magnetic flux through the coil is:


"\Phi_i = NBA"

where "N = 100", "B= 0.74T, A = 0.20m \u00d7 0.20 m=0.04m^2". The final flux is:


"\Phi_f = NBA\cos\theta"

where "\theta=60\degree". Thus, the change is:


"\Delta\Phi = \Phi_f - \Phi_i = NBA(\cos\theta -1)"

and the emf:


"\mathcal{E} = \dfrac{NBA(1-\cos\theta)}{\Delta t}\\
\mathcal{E} = \dfrac{100\cdot 0.74\cdot 0.04\cdot (1-\cos60\degree)}{10} = 0.148V"

Answer. 0.148V.